OpenStudy (anonymous):

Write in standard form an equation of the line passing through the given point with the given slope. slope = –7; (–2, –2) a. –7x + y = –16 b. 7x + y = 16 c. 7x – y = –16 d. 7x + y = –16

OpenStudy (anonymous):

Do you know the slope formula?

OpenStudy (anonymous):

y = Mx = b

OpenStudy (anonymous):

Y = Mx + b

OpenStudy (anonymous):

It's y=mx+b

OpenStudy (anonymous):

Now that you know the formula, where do you put the -7?

OpenStudy (anonymous):

im not sure?

OpenStudy (anonymous):

Standard form is Ax + By = C

OpenStudy (anonymous):

Do you have a math textbook with you?

OpenStudy (anonymous):

Not on me. i lent it to my friend

OpenStudy (anonymous):

http://www.mathwarehouse.com/algebra/linear_equation/standard-form-equation-of-a-line.php If you check out this website, it'll help you. If you still don't get it, you can tell me

OpenStudy (anonymous):

Im still confused?

OpenStudy (anonymous):

One second. Let me explain

OpenStudy (anonymous):

you have my attention

OpenStudy (anonymous):

y = mx + b is an equation. The question gives you the slope, and a pair of coordinates (-2,-2) and a slope of -7. The first -2 in the coordinate stands for x. The second -2 in the coordinate stands for y. M is the slope, and -7 stands for the M. Plug those numbers into the equation..

OpenStudy (anonymous):

and you get.. -2 = -7(-2) + b

OpenStudy (anonymous):

Solve that for me

OpenStudy (anonymous):

-2 = 14 +b from there im stuck again?

OpenStudy (anonymous):

Right. You need to subtract the 14 on both sides

OpenStudy (anonymous):

-16 = B

OpenStudy (anonymous):

Right! Now plug it into y=mx+b y=-7x-16

OpenStudy (anonymous):

Rearrange that into standard form.

OpenStudy (anonymous):

A is the only reasonable answer. -7x + y = -16

OpenStudy (anonymous):

-7x + y = -16

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

Do you understand it now?

OpenStudy (anonymous):

Yeah. i got lost on solving for the slope intercept form. but i get it now

OpenStudy (anonymous):

You going to close the question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

taking Notes